Ashley is 20 years older than Jacob. Five tears ago, she was eleven times as old as he was. How old are they now? Show equation and check.
Answer by stanbon(75887)

Ashley is 20 years older than Jacob. Five tears ago, she was eleven times as old as he was. How old are they now?
Equations:
a = j + 20
a-5 = 11(j-5)
----
Substitute for "a" and solve for "j":
j+20-5 = 11j-55
10j = 50
j = 5 yrs (Jason's age now)
----
Solve for "a":
a = j+20
Ashley age = 25 yrs
=======================
Cheers,
Stan H.